OpenStudy (anonymous):

What is two-fifths divided by six-sevenths? twelve thirty-fifths two and eleven-twelfths seven-fifteenths two and one-seventh

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

the rule for dividing fractions is flip the 2nd fraction and multiply so \[\frac{2}{5} \div \frac{6}{7}....becomes.....\frac{2}{5} \times \frac{7}{6} = \]

OpenStudy (anonymous):

When you divide fractions, you must use the reciprical of the fraction after the division sign and multiply. 2/5 divided by 6/7 becomes 2/5 * 7/6 = 14/30 = 7 /15